### v2

@@ 4.3. Robustness Evaluation
Order robustness. We evaluate the robustness of Oreo to variations in the order of retrieved documents by shuffling the top-5 retrieved results and comparing its performance against the original document order. The results for five datasets are presented in Table 5. From the table we can see that, Oreo consistently maintain the performance on five datasets (with ±$\pm 0.003$ to ±$\pm 0.027$). This highlights that Oreo is order- or position-agnostic. Even the retrieved chunks are suboptimally ranked or presented in an arbitrary order, Oreo can still effectively capture and synthesize essential information as long as the evidence exists in the chunks. This capability is largely attributed to Oreo ’s inherent reordering feature during context reconstruction, enabling it to function as an implicit reranker. Such robustness is particularly valuable for mitigating the ”lost-in-the-middle” (Liu et al., 2023) phenomenon, where the order of relevant information may influence the downstream generator’s performance.
4.4. Generalizability Evaluation
To evaluate the cross-dataset generalizability of Oreo , we assess its transferring capability by applying models trained on one dataset to tasks in a different dataset without any fine-tuning. This approach tests Oreo ’s ability to generalize its context reconstruction and synthesis capabilities to unseen query distributions. Specifically, we examine performance when using a model trained on PopQA to generate answers for NQ and a model trained on 2QWA to process HQA queries. The results, presented in Table 6, reveal that Oreo achieves competitive performance in zero-shot setting. For example, the model trained on PopQA achieves a score of 0.4352 when applied to NQ, only slightly lower than the performance of being specific trained (i.e 0.4413 and 0.4682). Similarly, using the 2WQA-trained model on HQA yields a score of 0.6344, closely matching the intra-dataset score of 0.6384. These findings demonstrate Oreo ’s ability to generalize its context reconstruction to similar QA types effectively, even under query distribution shifts. Its strong performance across datasets highlights its robustness and adaptability, making it a promising solution for open-domain QA tasks that require flexibility in handling diverse knowledge sources and query structures.
@@ 5.1. Post-retrieval Enhancement for RAG
Reranking. Rerankers reorder and prioritize retrieved documents to emphasize the most relevant results. They typically operate sequentially or iteratively after retrieval, leveraging various criteria such as semantic relevance between query and passages (Glass et al., 2022; Hofstätter et al., 2023), connections among documents (Dong et al., 2024), the majority of reader predictions (Mao et al., 2021), and utility for generation (Ma et al., 2023a). Rerankers are usually based on cross-encoder (e.g BGE (Xiao et al., 2023), Mixedbread (Li and Li, 2023)), multi-vector models (e.g ColBert (Khattab and Zaharia, 2020; Santhanam et al., 2022)). Recent works explore using LMs as rerankers (e.g, RankT5 (Zhuang et al., 2023), RankZephyr (Pradeep et al., 2023), RankGPT (Sun et al., 2023b)).
Filtering. Filtering is the process of selectively removing irrelevant, redundant, or low-quality information from retrieved documents to enhance their usefulness. Filtering can be performed at different granularity levels such as passage-level (Asai et al., 2022; Yoran et al., 2023; Asai et al., 2023), sentence-level (Wang et al., 2023a; Hwang et al., 2024) or finer token-level (Jin et al., 2024; Anderson et al., 2022; Jiang et al., 2024).
@@
Post verification and correction. Some studies incorporate post evaluation (Yan et al., 2024), fact-checking (Liu et al., [n. d.]) or attribution (Gao et al., 2023a) mechanisms to further solidify the accuracy and reliability of the retrieved documents and responses. For example, (Yan et al., 2024) uses a retrieval evaluator accessing the relevance of the retrieved documents and estimating a correctness confidence degree, then self-correct the retrieved results. RETA-LLM (Liu et al., [n. d.]) includes a fact checking module to verify whether there exist factual errors in the generated answers. by using a natural language understanding (NLU) model. CHAIN-OF-NOTE (CON) (Yu et al., 2024) accesses relevance of each retrieved document to the query by generating sequential reading notes. These notes are integrated during final answer generation.
Summarizing or compressing. Summarization and compression of retrieved content aim to create concise, faithful contexts that enhance the efficiency and effectiveness of downstream generation (Xu et al., 2024b). Several methods have been proposed to achieve this. (Xu et al., 2024b) proposed dual-encoder compressors for extractive and abstractive to eliminate irrelevant context and summarize upon the remaining content. Selective-Context (Li et al., 2023) employs a small LM compute the self-information of each lexical uni and then drop the less informative content. LongLLMLingua (Jiang et al., 2024) proposes coarse-grained compression via document-level perplexity and fine-grained compression of remaining text based on token perplexity.
5.2. Reinforcement Learning for Large Language Models
Reinforcement learning for Language Models (RL4LM) has emerged as a transformative technique to further enhance LLMs’ performance during the post-training process (Cao et al., 2024b; Pternea et al., 2024). Traditional RL4LM usually involves a reward model, for example using PPO (Schulman et al., 2015) to update the policy model (e.g InstructGPT (Ouyang et al., 2022), GPT-4 (Achiam et al., 2023)). Some RL4LM such as Direct Preference Optimization (DPO) (Rafailov et al., 2024) and Reward-aware Preference Optimization (RPO) (Adler et al., 2024) get rid of reward model to provide more stable and computationally efficient solutions (e.g Qwen 2 (Chu et al., 2024) and Nemotron-4 (Adler et al., 2024)). Common goals of RL4LM include improving performance of downstream NLP tasks (Deng et al., 2022; Ghalandari et al., 2022; Ramamurthy et al., 2023), minimizing data and resource dependencies (Zhang et al., 2022), aligning model outputs with user intent, values and goals (Ouyang et al., 2022), and adhering to responsible AI principles (Bai et al., 2022a, b). Human feedback can be integrated into the framework by constructing preference datasets, which are then used to fine-tune both the policy and reward models (also termed as Reinforcement Learning from Human Feedback (RLHF)) (Bai et al., 2022a; Ouyang et al., 2022; Hu et al., 2023). Some studies also explore RL4LM without human feedback (Rafailov et al., 2024) or replaced with AI feedback (Bai et al., 2022b; Yuan et al., 2024) by distillation from LLMs (Cui et al., 2023; Park et al., 2024), prompting LLMs as reward functions (Kwon et al., 2023; Lee et al., 2023; Zhang et al., 2024a), and self-rewarding (Yuan et al., 2024), or using performance-based metrics such as fluency or coherence (Ghalandari et al., 2022), and task-specific constraints over the distribution of language (Ramamurthy et al., 2023; Wu et al., 2024b). In the specific domain of RAG, RRAML (Bacciu et al., 2023) employs RL to train an efficient retriever to search for relevant information in in arbitrarily large databases. PRCA (Yang et al., 2023) applies RL to further fine-tune the context extracted by a model pre-trained on domain-specific abstractive summarization tasks to optimize the reward for the generator. BIDER (Jin et al., 2024) adopts RL to bridge the inconsistency between the retriever and generator.
6. Conclusion
We have presented Oreo - a lightweight and pluggable module designed to enhance the performance of RAG systems by reconstructing retrieved document chunks and mitigating the potential knowledge inconsistencies between the retriever and generator. Upon receiving document chunks from the retriever, Oreo efficiently filters out irrelevant, redundant, and distracting content, transforming them into a concise and query-supportive context. These reconstructed contexts effectively guide the generator toward producing accurate answers for open-domain QA tasks. Notably, Oreo can be seamlessly integrated with arbitrary retrievers, generators, or other RAG components without requiring significant adjustments or modifications. Experimental results demonstrate Oreo ’s effectiveness in downstream tasks, its efficiency in compressing context while improving performance, and its robustness in handling noisy and imperfectly ranked document chunks.
Limitations and future works. Despite Oreo ’s demonstrated effectiveness in short-form factual QA tasks, certain limitations remain. First, to achieve a high compression rate, Oreo enforces a maximum token count during context synthesis. While effective for most factual QA tasks, this constraint may inadvertently exclude vital information, particularly in complex tasks that require multi-hop reasoning or synthesizing extensive, interconnected evidence. Future work should investigate the trade-off between token counts and different tasks, and explore dynamic token allocation strategies that adapt to the query type (e.g multi-hop QA, long-form QA) and consider latency requirements during inference. Second, while Oreo has shown resilience to perturbed document orders and noise-injected context, it has not been thoroughly tested in adversarial environments, such as those involving misinformation or conflicting evidence in retrieved documents. Future research should consider evaluating and improving Oreo ’s robustness in such challenging scenarios. Third, the current evaluation of Oreo relies heavily on the performance of downstream tasks, which serve as indirect assessments and may introduce bias. Developing advanced evaluation schemes (e.g LLM-as-a-jedge (Gu et al., 2024)), will be essential for providing a more direct and reliable assessment of the quality of Oreo ’s reconstructions. Addressing these challenges could unlock the full potential of Oreo in more diverse and demanding real-world applications.
7. Appendices
Table 7. Prompt template used for data collection.
Prompt Input: Your task is to decompose the question, extract and abstract supporting information from the context to answer the question. Your output should mention all entities involved in the question, supporting sentences and rationals to all sub-questions from the context. If the conetxt doesn’t provide information to answer the question, output ’[UNKNOWN]’. Output the ¡Output¿ part only.
Example1:
¡Question¿: Where was the director of film The Fascist born?
¡Context¿: {Retrieved document chunks}
¡Output¿: Luciano Salce, the director of the satirical film The Fascist, was born on September 25, 1922, in Rome, Italy. Salce was an Italian filmmaker, actor, and screenwriter known for his ability to blend comedy with social and political critique.
Example2:
¡Question¿: what is the number 1 sport in the usa?
¡Context¿: {Retrieved document chunks}
¡Output¿: American football is the most popular sport in the United States followed by basketball, baseball, and soccer.
Example3:
¡Question¿: What was the first English monastery to be sacked by the Norsemen?
¡Context¿: {Retrieved document chunks}
¡Output¿: Vikings attacked the monastery at Lindisfarne on June 8, 793, which is the first recorded Viking raid on an English monastery.
Example4:
¡Question¿: Kate Philips played which wife of Henry VIII in ’Wolf Hall’?
¡Context¿: {Retrieved document chunks}
¡Output¿: Kate Phillips played Abigail Williams in ”The Crucible” at the West Yorkshire Playhouse, and then went on to film her scenes for the BBC’s adaptation of ”Wolf Hall” in which she played Jane Seymour, Henry VIII’s third wife.
Example5:
¡Question¿: Lokomotiv Yaroslavl was the team founded in 2011 after the plane crash near which airport?
¡Context¿: {Retrieved document chunks}
¡Output¿: Lokomotiv Yaroslavl Hockey Club Lokomotiv, also known as Lokomotiv Yaroslavl, is a Russian professional ice hockey team. On 7 September 2011, nearly the entire team perished in the Lokomotiv Yaroslavl plane crash. The aircraft ran off the runway before lifting off, struck a tower mast, caught fire and crashed from the end of the runway of Tunoshna Airport on the Volga River bank.
{Question}
{Retrieved document chunks}
Output: {Output}
Table 8. Bootstraping Prompt Template
Prompt Input: You are given a question, a set of document chunks, a correct answer, extract evidences and supporting information from the chunks and generate rationales how these information derive the correct answer.
Example1:
¡Question¿: What nationality were social anthropologists Alfred Gell and Edmund Leach?
¡Chunks¿: {Retrieved document chunks}
¡Correct answer¿: British.
¡Output¿: Both Alfred Gell and Edmund Leach were British. They were educated and primarily worked within the United Kingdom’s academic framework. Their national and professional affiliations firmly establish their British nationality.
Example2:
¡Question¿: Crucible is a geodemography computer system created by a company that has stores in how many countries?
¡Chunks¿: {Retrieved document chunks}
¡Correct answer¿: 12.
¡Output¿: Crucible is a geodemography computer system created by Tesco, a multinational grocery and general merchandise retailer. Tesco has stores in 12 countries as of recent data, so 12 is the answer.
Example3:
¡Question¿: What word is in both the genre of Muhammed Suiçmez’s band and the genre of Dave Meniketti’s band?
¡Chunks¿: {Retrieved document chunks}
¡Correct answer¿: Metal.
¡Output¿: Necrophagist is known for its death metal style. Y&T is often classified under the broader category of heavy metal. So the answer is mental.
{Question}
{Retrieved document chunks}
{Correct answer}
Output: {output}

### v3

@@ 4.3. Robustness Evaluation
Figure 5. Left (a) - Comparison of number of input tokens for generator and QA performance across different context types. Right (b) - Comparison of end-to-end inference time (measured in seconds) by using different types of context.
[…]
tokenvsperf
Order robustness. We evaluate the robustness of Oreo to variations in the order of retrieved documents by shuffling the top-5 retrieved results and comparing its performance against the original document order. The results for five datasets are presented in Table 4. From the table we can see that, Oreo consistently maintain the performance on five datasets (with ±$\pm 0.003$ to ±$\pm 0.027$). This highlights that Oreo is order- or position-agnostic. Even the retrieved chunks are suboptimally ranked or presented in an arbitrary order, Oreo can still effectively capture and synthesize essential information as long as the evidence exists in the chunks. This capability is largely attributed to Oreo ’s inherent reordering feature during context reconstruction, enabling it to function as an implicit reranker. Such robustness is particularly valuable for mitigating the ”lost-in-the-middle” (Liu et al., 2023b) phenomenon, where the order of relevant information may influence the downstream generator’s performance.
4.4. Generalizability Evaluation
To evaluate the cross-dataset generalizability of Oreo , we assess its transferring capability by applying models trained on one dataset to tasks in a different dataset without any fine-tuning. This approach tests Oreo ’s ability to generalize its context reconstruction and synthesis capabilities to unseen query distributions. Specifically, we examine performance when using a model trained on PopQA to generate answers for NQ and a model trained on 2WQA to process HotpotQA queries. We report the detailed results in Table 9 in Appendix E, which demonstrate that Oreo achieves competitive performance in the zero-shot setting. For example, the model trained on PopQA achieves a score of 0.4352 when applied to NQ, only slightly lower than the performance of being specific trained (ie. 0.4413 and 0.4682). Similarly, using the 2WQA-trained model on HotpotQA yields a score of 0.6344, closely matching the intra-dataset score of 0.6384. These findings demonstrate Oreo ’s ability to generalize its context reconstruction to similar QA types effectively, even under query distribution shifts. Its strong performance across datasets highlights its robustness and adaptability, making it a promising solution for open-domain QA tasks that require flexibility in handling diverse knowledge sources and query structures.
@@ 4.4. Generalizability Evaluation
4.5. Faithfulness and Completeness Evaluation
Apart from the downstream task performance, the quality of context generated by Oreo is essential, esp. the factual accuracy (faithfulness) and coverage of relevant information (completeness) with respect to the original retrieved passages. To this end, we conduct an evaluation of both faithfulness and completeness to ensure that Oreo produces context that is not only concise but also reliable and fully representative of the source passages.
We adopt the LLM-as-a-judge framework (Gu et al., 2024) to systematically assess these dimensions. In particular, we prompt Qwen2.5-Instruct (Yang et al., 2024) to evaluate the generated context by assigning faithfulness and completeness scores on a 0–10 scale. Faithfulness reflects the degree to which the context remains factually grounded in the original retrieved content, avoiding hallucinations or the introduction of extraneous information. Completeness assesses whether the context sufficiently captures all salient and relevant details from the original passages with respect to the query. To promote transparency and interpretability, the model is also asked to provide a rationale supporting each score. Prompts for scores and explanation generation are provided in Appendix D.
We provide evaluation results in Figure 7 in Appendix D.2 Among all methods, Oreo consistently achieves the highest scores across all datasets, excelling in both completeness and faithfulness, with standout scores on complex datasets like HotpotQA (8.87 completeness, 8.97 faithfulness). CompAct emerges as a strong second-best, showing strong balance and high faithfulness, especially on HotpotQA and 2WQA. Refiner delivers moderately competitive results, generally maintaining factual consistency but showing limitations in coverage. EXIT demonstrates lower overall performance, especially struggling on more demanding datasets such as 2WQA. In contrast, LLMLingua and LLMLingua-2 produce the weakest results, with both completeness and faithfulness significantly lower across all datasets, likely due to aggressive filtering or compression strategies that sacrifice critical information.
4.6. Ablation Study
We conduct an ablation study to investigate the effect of varying context lengths generated by Oreo . Specifically, we progressively increase the minimum token threshold for context generation from 30 to 300 tokens, in increments of 30, while fixing the number of retrieved passages to top-5. The results of downstream task performance are summarized in Figure 8 in Appendix F. Our findings indicate a consistent performance improvement across all datasets as the minimum context length increases, with gains being more pronounced in the earlier stages (from 30 to 180 tokens). These improvements suggest that extending the context allows the model to access a broader set of relevant evidence, improving its ability to synthesize accurate responses. However, beyond a threshold, typically between 240 and 270 tokens, we observe a performance plateau or marginal decline. This indicates diminishing returns with excessively long contexts. While longer windows can potentially capture more relevant details, they also risk introducing extraneous, redundant, or weakly relevant content, which can dilute the core information necessary for accurate answers.
@@ 5.1. Post-retrieval Enhancement for RAG
Reranking. Rerankers reorder and prioritize retrieved documents to emphasize the most relevant results. They typically operate sequentially or iteratively after retrieval, leveraging various criteria such as semantic relevance between query and passages (Glass et al., 2022; Hofstätter et al., 2023), connections among documents (Dong et al., 2024a), the majority of reader predictions (Mao et al., 2021), and utility for generation (Ma et al., 2023a). Rerankers are usually based on cross-encoder (eg., BGE (Xiao et al., 2023), Mixedbread (Li and Li, 2023)), multi-vector models (eg., ColBert (Khattab and Zaharia, 2020; Santhanam et al., 2022)). Recent works also explore using LMs as rerankers (eg., RankT5 (Zhuang et al., 2023), RankZephyr (Pradeep et al., 2023), RankGPT (Sun et al., 2023b), DPA-RAG (Dong et al., 2024b)).
Post verification and correction. Some studies incorporate post-hoc evaluations to improve factuality and relevance of retrieved documents. Examples include relevance evaluators (Yan et al., 2024), fact-checkers (Liu et al., [n. d.]), attribution (Gao et al., 2023a; Yu et al., 2024) and multi-agent (Xu et al., 2024a) mechanisms to further solidify the accuracy and reliability of the retrieved documents and responses.
Compressing. Compression methods condense retrieved content to improve efficiency and focus. These methods can be broadly categorized into lexical-based and embedding-based approaches. Lexical-based methods usually involve summarization techniques (Xu et al., 2024c; Liu et al., 2023a) to retain essential information, semantic filtering to remove low-importance tokens, and both extractive and abstractive strategies for eliminating irrelevant context (Xu et al., 2024c). Some approaches compute the self-information of lexical units to discard less informative content (Li et al., 2023), or apply token-level filtering based on perplexity (Jiang et al., 2024). Embedding-based methods, on the other hand, condense documents into fixed-size representations in the embedding space, recent works include xRAG (Cheng et al., 2024) and PISCO (Louis et al., 2025). Our work falls falls within the lexical-based compression group.
5.2. Reinforcement Learning for Large Language Models
Reinforcement learning for Language Models (RL4LM) has emerged as a transformative technique to further enhance LLMs’ performance during the post-training process (Cao et al., 2024b; Pternea et al., 2024). Traditional RL4LM usually involves a reward model, for example using PPO (Schulman et al., 2015) to update the policy model (eg., InstructGPT (Ouyang et al., 2022), GPT-4 (Achiam et al., 2023)). Some RL4LM such as Direct Preference Optimization (DPO) (Rafailov et al., 2024) and Reward-aware Preference Optimization (RPO) (Adler et al., 2024) get rid of the reward model to provide more stable and computationally efficient solutions (eg., Qwen 2 (Chu et al., 2024) and Nemotron-4 (Adler et al., 2024)). Common goals of RL4LM include improving performance of downstream NLP tasks (Deng et al., 2022; Ghalandari et al., 2022; Ramamurthy et al., 2023), minimizing data and resource dependencies (Zhang et al., 2022), aligning model outputs with user intent, values and goals (Ouyang et al., 2022), and adhering to responsible AI principles (Bai et al., 2022a, b). Human feedback can be integrated into the framework by constructing preference datasets, which are then used to fine-tune both the policy and reward models (also termed as Reinforcement Learning from Human Feedback (RLHF)) (Bai et al., 2022a; Ouyang et al., 2022; Hu et al., 2023). Some studies also explore RL4LM without human feedback (Rafailov et al., 2024) or replaced with AI feedback (Bai et al., 2022b; Yuan et al., 2024) by distillation from LLMs (Cui et al., 2023; Park et al., 2024), prompting LLMs as reward functions (Kwon et al., 2023; Lee et al., 2023; Zhang et al., 2024a), and self-rewarding (Yuan et al., 2024), or using performance-based metrics such as fluency or coherence (Ghalandari et al., 2022), and task-specific constraints over the distribution of language (Ramamurthy et al., 2023; Wu et al., 2024b). In the specific domain of RAG, RRAML (Bacciu et al., 2023) employs RL to train a retriever in arbitrarily large databases. PRCA (Yang et al., 2023) applies RL to fine-tune the context to optimize the reward for the generator. BIDER (Jin et al., 2024) adopts RL to bridge the inconsistency between the retriever and generator.
6. Conclusion
We have presented Oreo - a lightweight and pluggable module designed to enhance the performance of RAG systems by reconstructing retrieved document chunks and mitigating the potential knowledge inconsistencies between the retriever and generator. Upon receiving document chunks from the retriever, Oreo efficiently filters out irrelevant, redundant and distracting content, transforming them into a concise and query-supportive context. These reconstructed contexts effectively guide the generator toward producing accurate answers for open-domain QA tasks. Notably, Oreo can be seamlessly integrated with arbitrary retrievers, generators, or other RAG components without requiring significant adjustments or modifications. Experimental results demonstrate Oreo ’s effectiveness in downstream tasks, its efficiency in compressing context while improving performance, and its robustness in handling noisy and imperfectly ranked document chunks.
Limitations. While Oreo shows strong performance on open-domain QA tasks, it has some limitations. First, its aggressive compression may omit essential information in complex settings like multi-hop or long-form QA. Second, Oreo has not been systematically tested in adversarial retrieval scenarios (Wang et al., 2025) involving conflicting or deceptive content. Third, current evaluations rely heavily on indirect metrics such as downstream QA accuracy or LLM judgments, which may introduce bias.
Future Work. Future research will explore adaptive compression strategies that dynamically allocate token budgets based on query complexity and task type. Robustness to adversarial or noisy retrieval scenarios also warrants closer investigation, especially for high-stakes domains. We are also interested in developing more principled and fine-grained evaluation frameworks to better understand the trade-offs between compression, informativeness and faithfulness in context reconstruction. Lastly, to address sparsity in rewards, a promising direction for future work is to develop progress-based RL frameworks that incorporate intermediate quality assessments of Oreo ’s reconstructed context, providing denser and more fine-grained rewards to enable more stable and efficient policy learning.
Appendix A Prompt Templates for Data Collection
A.1. Prompt Templates for Data Collection
Input: Your task is to decompose the question, extract and abstract supporting information from the context to answer the question. Your output should mention all entities involved in the question, supporting sentences and rationals to all sub-questions from the context. If the conetxt doesn’t provide information to answer the question, output ’[UNKNOWN]’. Output the ¡Output¿ part only. Example1: ¡Question¿: Where was the director of film The Fascist born? ¡Context¿: {Retrieved document chunks} ¡Output¿: Luciano Salce, the director of the satirical film The Fascist, was born on September 25, 1922, in Rome, Italy. Salce was an Italian filmmaker, actor, and screenwriter known for his ability to blend comedy with social and political critique. Example2: ¡Question¿: what is the number 1 sport in the usa? ¡Context¿: {Retrieved document chunks} ¡Output¿: American football is the most popular sport in the United States followed by basketball, baseball, and soccer. Example3: ¡Question¿: What was the first English monastery to be sacked by the Norsemen? ¡Context¿: {Retrieved document chunks} ¡Output¿: Vikings attacked the monastery at Lindisfarne on June 8, 793, which is the first recorded Viking raid on an English monastery. Example4: ¡Question¿: Kate Philips played which wife of Henry VIII in ’Wolf Hall’? ¡Context¿: {Retrieved document chunks} ¡Output¿: Kate Phillips played Abigail Williams in ”The Crucible” at the West Yorkshire Playhouse, and then went on to film her scenes for the BBC’s adaptation of ”Wolf Hall” in which she played Jane Seymour, Henry VIII’s third wife. Example5: ¡Question¿: Lokomotiv Yaroslavl was the team founded in 2011 after the plane crash near which airport? ¡Context¿: {Retrieved document chunks} ¡Output¿: Lokomotiv Yaroslavl Hockey Club Lokomotiv, also known as Lokomotiv Yaroslavl, is a Russian professional ice hockey team. On 7 September 2011, nearly the entire team perished in the Lokomotiv Yaroslavl plane crash. The aircraft ran off the runway before lifting off, struck a tower mast, caught fire and crashed from the end of the runway of Tunoshna Airport on the Volga River bank. {Question} {Retrieved document chunks} Output: {Output}
A.2. Prompt Template for Boostraping Data Generation
Input: You are given a question, a set of document chunks, a correct answer, extract evidences and supporting information from the chunks and generate rationales how these information derive the correct answer. Example1: ¡Question¿: What nationality were social anthropologists Alfred Gell and Edmund Leach? ¡Chunks¿: {Retrieved document chunks} ¡Correct answer¿: British. ¡Output¿: Both Alfred Gell and Edmund Leach were British. They were educated and primarily worked within the United Kingdom’s academic framework. Their national and professional affiliations firmly establish their British nationality. Example2: ¡Question¿: Crucible is a geodemography computer system created by a company that has stores in how many countries? ¡Chunks¿: {Retrieved document chunks} ¡Correct answer¿: 12. ¡Output¿: Crucible is a geodemography computer system created by Tesco, a multinational grocery and general merchandise retailer. Tesco has stores in 12 countries as of recent data, so 12 is the answer. Example3: ¡Question¿: What word is in both the genre of Muhammed Suiçmez’s band and the genre of Dave Meniketti’s band? ¡Chunks¿: {Retrieved document chunks} ¡Correct answer¿: Metal. ¡Output¿: Necrophagist is known for its death metal style. Y&T is often classified under the broader category of heavy metal. So the answer is mental. {Question} {Retrieved document chunks} {Correct answer} Output: {output}
Appendix B Statistics and Experimental Setups for Datasets
Table 5. Dataset statistics, retrievers and evaluation metrics. EM -Exact Match, F1 - Unigram F1
Dataset # Train (k) # Test (k) Retriever Precision@5 Task Metric
PopQA 6.5 1.4 Contriver 0.287 Extractive single-hop QA EM
NaturalQuestions 28.3 3.6 DPR 0.33 Extractive single-hop QA EM
TriviaQA 30.1 11.3 Contriver 0.43 Extractive single-hop QA EM
HotpotQA 20.7 5.6 Contriver 0.137 Abstractive multi-hop QA F1
2WikiMultiHopQA 20.7 12.6 BM25 0.07 Abstractive multi-hop QA F1
Appendix C Parameter Settings
We detail the key hyperparameters and configurations used across all experiments in Table 6. Specifically, CML and RL represents contrastive multitask learning and reinforcement learning respectively.
Parameter Value
$\eta$ (CML) 0.01
$\alpha$ (CML) 0.5
$\epsilon$ (RL) 0.2
$\gamma$, $\lambda$(RL) 0.95
Top-k (RL) 4
Top-p (RL) 0.95
Appendix D LLM-as-a-Judge For Faithfulness and Completeness Evaluation
D.1. Prompts for Qwen-2.5-Instruct
To directly evaluate the quality of context generated by Oreo , we employ Qwen-2.5-Instruct (Yang et al., 2024) as a reference model to assess two critical dimensions: faithfulness - how well the answer aligns with the retrieved passages, and completeness - to what extend does the generated context cover all essential information to correctly answer the query.
We design instructions for prompting Qwen-2.5-Instruct, as detailed in Table 7 and Table 8.
Table 7. Qwen-2.5-Instruct used for faithfulness evaluation
Faithfulness evaluation prompt Input: ”””You are an expert evaluator assessing the **faithfulness** of a generated context with respect to the original passages. You will be given a query, a set of original passages, and a generated context. Your task is to determine how accurately the generated context reflects the facts and meanings in the original passages, focusing only on the information required to answering the query. Consider whether the context includes hallucinated information, omits key facts, or misrepresents any content. Rate the faithfulness on a scale from 0 to 10: - 0: The generated context is entirely unfaithful or unrelated to the original passages. - 10: The generated context is fully faithful, with no hallucinations, distortions, or omissions of relevant information. Output your score (a float between 0 and 10), followed by a concise explanation of your reasoning. ”””
Text: {text}
Output: {Output}
Table 8. Qwen-2.5-Instruct used for completeness evaluation
Completeness evaluation prompt Input: ”””You are an expert evaluator assessing the **completeness** of a generated response. You will be given a query, a set of original passages, and a generated context intended to answer the query. Your task is to rate how thoroughly the generated context covers all necessary information from the original passages required to answer the query. The context should not omit relevant details, and should avoid adding any external or fabricated content. Rate the completeness on a scale from 0 to 10: - 0: The generated context provides no useful information for answering the query. - 10: The generated context includes all necessary information to fully and correctly answer the query. Output your score (a float between 0 and 10), followed by a brief explanation of your reasoning. ”””
Text: {text}
Output: {Output}
D.2. Scoring Results
Figure 7 presents the completeness and faithfulness scores evaluated by Qwen-2.5-Instruct, demonstrating that Oreo achieves the highest performance on both metrics.
Figure 7. Completeness and faithfulness evaluation by Qwen-2.5-Instruct
[…]
qweneval
Appendix E Generalizability Evaluation
To evaluate cross-dataset generalizability, we test Oreo ’s transferability by applying models trained on one dataset to a different one without fine-tuning. This assesses Oreo ’s ability to reconstruct and synthesize context under unseen query distributions. Specifically, we evaluate models trained on PopQA for NQ, and on 2WQA for HotpotQA. Detailed results are provided in Table 9.
@@ Appendix E Generalizability Evaluation
Appendix F Ablation Study on Generated Token Numbers
We perform an ablation study to assess how varying Oreo ’s context length affects downstream performance. By increasing the minimum token threshold from 30 to 300 (in steps of 30) while keeping the top-5 retrieved passages fixed, we observe performance trends summarized in Figure 8.
Figure 8. Performance of Oreo generating different lengths of context across five datasets
[…]
ablation